# Prompts As Programs: A Structure-Aware Approach to Efficient Compile-Time Prompt Optimization

### v1

@@ Abstract
Large language models (LLMs) can now handle longer and more complex inputs, which facilitate the use of more elaborate prompts. However, prompts often require some tuning to improve performance for deployment. Recent work has proposed automatic prompt optimization methods, but as prompt complexity and LLM strength increase, many prompt optimization techniques are no longer sufficient and a new approach is needed to optimize meta prompt programs. To address this, we introduce SAMMO, a framework for compile-time optimizations of metaprompt programs, which represent prompts as structured objects that allows for a rich set of transformations that can be searched over during optimization. We show that SAMMO generalizes previous methods and improves the performance of complex prompts on (1) instruction tuning, (2) RAG pipeline tuning, and (3) prompt compression, across several different LLMs. We make all code available open-source at https://github.com/microsoft/sammo.
@@ 1 Introduction
{forest}
for tree= rectangle,draw, font= [Metaprompt $\pi[X]$,fill=lightergray [RenderText()
$\theta_{\text{text}}$ = “In the following task,
you will have to classify…” ] [RenderSection(), [RenderText()
$\theta_{\text{text}}$
= “Here are
some examples:” ] [RenderData()
$\theta_{\text{format}}$ = “json”
$\theta_{\text{fewshot}}$ = [{“Good meal!”:
“positive”}, …], ] ] [RenderSection($X$),fill=lightergray [RenderText()
$\theta_{\text{text}}$
= “Classify this:”] [RenderData($X$)
$\theta_{\text{format}}$ = “json”,fill=lightergray ] ] ]
Figure 1: A simple metaprompt for a review classification task. Sammo represents metaprompt as a dynamic function graph where each node computes a new value based on the results of its children, the input data $X$ as well as node-specific parameters $\theta$. Here, nodes that use the input data $X$ are marked in light gray.
@@
With the recent development of large language models (LLMs) such as Mixtral 8x7B Jiang et al. (2024) or GPT-4, it is now possible to provide LLMs with longer inputs including richer context and more detailed instructions. As a result, the complexity of prompts has increased, including longer strings of instructions, demonstrations or examples, and specification of more structured output. These lengthy instructions are often reused as dynamic templates (aka metaprompts), where static information (eg, instructions) is combined with input-dependent information (e.g., user queries, retrieved documents in Retrieval Augmented Generation (RAG) systems) at runtime to generate the desired output. There is a clear trend towards prompts becoming complex programs in and of themselves.
@@
To achieve an acceptable level of accuracy for deployment, prompts are often manually fine-tuned. This process typically involves adding information to handle exceptions and corner cases. To improve this process, there has been a great deal of recent work on automatic optimization methods. The first wave of work in this area focused on simpler prompts (i.e., with less structure) and unstructured mutators such as text rewrite operations Chen et al. (2023); Pryzant et al. (2023); Zhou et al. (2023). The second wave of work has focused on optimizing prompts with more complex programmatic structure such as meta-prompts. Initial work in this direction focused on applying textual mutators to different prompt components Fernando et al. (2023); Ye et al. (2023). Subsequent work considered both textual mutation and hyperparameter selection Sclar et al. (2023). However, with the increasing strength of LLMs and increasing complexity of prompt structure, many prompt optimization techniques are no longer applicable and a new approach is needed that is able to optimize metaprompt programs.
To address this, we introduce Sammo, a general purpose framework for compile-time optimizations of prompt programs. Individual operations and parts of the prompt are represented through components in a call graph, in spirit similar to DSpy (Khattab et al., 2023). However, Sammo goes beyond this previous work by (i) allowing this call graph to be changed (ii) also representing the internal structure of prompts and (ii) generalizing “compile-time” optimizations to all prompt components (eg. textual content, hyperparameters). It considers metaprompts as programs—to allow for modular construction of complex prompts and facilitate compilation of more effective prompts.
Specifically, Sammo represents metaprompts as structured objects, which allows for a much richer set of transformations to be searched over during optimization. We formalize the optimization as a problem of searching over the structure of a complex metaprompt $\pi$, which may include many structured components such as: task description, guidelines, examples, and input/output format. Note, a metaprompt is designed to be combined dynamically with different input data $X$ at run time, i.e., $\operatorname{LLM}(\pi[X])=Y$.
Figure 1 shows a possible structured program for a review classification task. Treating prompts as programs allows us to (i) treat the increasingly complex nature of metaprompts where older methods are not longer applicable and (ii) have general purpose mutation prompt operators to define a meaningful search space. Drawing from approaches for neural architecture search, Sammo carries out a genetic search over a set of mutation operators that can change the structure and information contained in non-trivial ways.
@@
Sammo is a general framework for prompt optimization in a black-box setting where the practitioner can only sample from the output of LLMs, reflecting common API capabilities Sun et al. (2022). We show that Sammo encompasses several instruction optimization and compression techniques as special cases. We demonstrate the utility of Sammo in three scenarios: (1) instruction tuning, (2) retrieval augmented generation (RAG) pipeline tuning, and (3) prompt compression for annotation tasks. Our experiments show the Sammo generalizes previous methods and provides gains of 10-100% in instruction tuning, 26-133% in RAG tuning, and over 40% in prompt compression in performance. Moreover, we show that complex prompts need to be optimized separately for each LLM and that gains are more pronounced for weaker LLMs. Code is available at https://github.com/microsoft/sammo under an MIT license.
2 Problem Definition & Notation
Let $\pi$ refer to a metaprompt function that takes input data $X$ and maps it to a string that is fed to an LLM to generate an output, i.e., $\hat{Y}=LLM(\pi[X])$. Our goal is to optimize the performance of $\pi$ by transforming it into a more performant metaprompt $\pi^{*}$ via modifications to its structure, parameters, and content. More precisely, our goal is to find a metaprompt $\pi$ with minimal loss across the entire data distribution:
@@ 2 Problem Definition & Notation
Here, $P(X,Y)$ is the distribution that the labeled samples $D_{s}$ are drawn from, $\Pi$ represents the set of all metaprompts, and $S_{\phi}\in\mathbb{R}$ is a scalarized loss function that specifies trade-offs between potentially multiple individual quality objectives $S_{i}$, (e.g., $S_{i}:=\mathcal{L}_{S_{i}}(Y,LLM(\pi[X]))$.
Since the data distribution $P(X,Y)$ is unknown and the evaluation cost is prohibitive, we resort to using the following sampled score in the spirit of empirical risk minimization (ERM):
@@
2.1 Compile-time Optimization
In this paper, we focus on optimization that can only be done once before deploying the metaprompt, which we refer to compile-time optimization and correponds to the problem of solving Eq. 2. More formally, the optimization itself is a function returning another function (the metaprompt):
$\hat{\pi}=\rho_{\text{compile}}(\Pi,D_{\text{train}},S_{\phi}).$ (3)
In this case, the optimization can only be done once before deploying the metaprompt. This is different from run-time optimization where the optimizer $\rho_{\text{run}}$ gets called with the filled in metaprompt, i.e., $\rho_{\text{run}}(\pi[X])=\tilde{X}$ and then the LLM is prompted as with $LLM(\tilde{X})$. This optimization procedure would need to be executed repeatedly for each different input data $X$. Since we aim to amortize the optimization cost over multiple uses of the metaprompt, we do not consider run-time optimization further. However, we note that run-time optimizations can compliment compile-time optimizations if sufficient resources are available.
Motivated by real-world needs, we make the following assumptions:
  1. 1.
    We only have small amounts of data for optimization (on the order of hundreds of examples). This represents a reasonable amount of data that can be hand-labeled; increasing the amount would limit the applicability in practice substantially.
  2. 2.
    The language model is a closed black box and does not output probabilities, reflecting recent changes in how access to LLMs is provided. However, we can sample from it: $\mathbf{y}\sim\operatorname{LLM}(Y\mid\pi[X])$.
2.2 Metaprompts As Programs
To search efficiently over the exponentially large space of all metaprompts $\Pi$, we consider their complex programmatic structure, which typically comprises parameterized components executed in sequence. More specifically, we represent the structure of $\pi$ as a directed acyclic function graph $G$, i.e., $\pi[X]=f_{G_{\pi}}(X)$. Here $G_{\pi}=(V,E)$ is a DAG with a single root node $v_{r}$. Directed edges $e_{ij}\in E$ represent parent ($v_{i}$) to child ($v_{j}$) relationships in the metaprompt structural components.
Each node $v\in V$ has a function type $\psi_{v}$ and static parameters $\theta_{v}$. We use $\theta_{v}$ to refer to both static textual input in the meta prompt that is common across calls (e.g., instructions) and any hyperparameters of the components (e.g., format specifiers). Each node also gets dynamic information from (1) $X$ (which may be transformed by its parents) and (2) messages from its children.
Then the graph is recursively evaluated from the rode node $v_{r}$ as follows:
$f_{G_{\pi}}(X)=\psi_{v_{r}}\bigg{(}(X\cup\psi_{\text{children}}^{r}),\theta_{v_{r}}\bigg{)},$
where
$\psi_{\text{\text{children}}}^{r}=\Big{\{}\psi_{v_{i}}\left(\dot{X}_{\psi_{v_{r}}}(X),(\theta_{v_{i}}\cup\theta_{v_{r}})\right)\Big{\}}_{v_{i}\>}$
$s.t.\>e_{ri}\in G_{\pi}.$
Here $X$ to refer to dynamic input that changes based on the call to the meta prompt (e.g., user query) and we note that $X$ may consist of a batch of more than one query. We use $\dot{X}_{\psi_{v}}(.)$ to refer to a possible transformation of the dynamic input that a parent $v$ can send its children, e.g., to perform minibatching.Note that the input can also be passed on directly without transformation, e.g., $\dot{X}_{\psi_{v}}(X)=X$.
We denote the space of metaprompts as $\Pi$, and each $\pi\in\Pi$ consists of a different program structure $G$. With this notation, we can see how searching over $\Pi$ in Eq. 2 amounts to combinatorial search over $G$. We initialize the search with a baseline prompt $\pi_{0}$ and generate successor functions through mutations to $G_{\pi_{0}}$. Mutators (see Sec. 3.1) operate on $G_{\pi}$ at the node or neighborhood level. They can reorder, replace, or delete nodes. They can also modify $\theta_{v}$ associated with a node (eg. by modifying instructions or parameter values).
We note the similarity of metaprompt search with neural architectures search—nodes functions that ignore $X$ can be seen as global parameters, $\psi_{v_{r}}$ correspond to projection operators or activation layers as they aggregate input from their predecessors.
To give some intuition for what this looks like in practice, Figure 1 shows a metaprompt for a review classification task which has already been instantiated with some input data $X$. It has a section with task instructions at the beginning, followed by a set of static in-context examples (although they could be selected dynamically as well) and the unlabeled input data. Note that the data format is explicitly represented—here it serializes data to a JSON array with each example being a dictionary.
@@ 3 Sammo: Structure-Aware Multi-objective Metaprompt Optimization
Figure 2: Sammo is a flexible framework for structured prompt optimization offers two classes of search algorithms, depending on the set of mutators used.
@@
11:     $\Pi_{\mathcal{C}}\!\leftarrow\!\!\Pi_{\mathcal{C}}\cup\textsc{Prune}(\Pi_{\text{nextgen}},D_{\text{train}},S_{\phi})$
@@
We outline Sammo, which is a general framework for optimizing the performance of metaprompts. Sammo uses general-purpose search algorithms with a rich set of mutation operators to explore the space of prompts. Figure 2 presents an overview of Sammo’s main components that allow for efficient prompt optimization. First, there is a base layer of programming primitives from which more complex prompt programs are being built (cf. Figure 1). These comprise the nodes of the metaprompt function graph $G$, which we described in Section 2.2. On top of that sits a layer of prompt mutators which are then used with the search algorithms in the top-most layer. We outline these two layers next.
@@ 3.1 Prompt Mutation Operators
Table 1: Examples for mutation operators, grouped by what part of a metaprompt $\pi$ they affect. Sammo allows for a rich set of operations whereas traditional prompt optimization techniques have only focused on operations that change the text.
@@
At the heart of Sammo optimization are mutation operators. Formally, a mutation operator is a probabilistic function $m:\Pi\times\Pi_{\Omega}\longrightarrow[0,1]$ that specifies how to transition from a metaprompt $\pi$ to an edited version $\pi^{\prime}\in\Pi_{\Omega}$. This structure-aware component of Sammo opens up a new class of operators, for example operators that only modify specific sections or paragraphs. These can range from trivial (e.g., rephrasing a sentence) to complex (e.g., inducing a new set of task instructions). We call an operator finite, if for all inputs $\pi$, the set of possible outputs is finite, and infinite otherwise. A mutation operator’s type will also determine what search algorithms can be used.
Table 1 shows a non-comprehensive set of mutation operators, grouped by what part of a metaprompt they change. Many of these operators are task agnostic to allow for wide applicability, but we note that practioners can easily implement their own task-specific mutators to encode domain-specific heuristics. To the best of our knowledge, Sammo is the first optimization method that can also optimize for large structural changes and data formatting.
3.2 Search Algorithms
There are two classes of search algorithms that Sammo provides. First, if all mutators used are finite, then Sammo can explore the space via enumerative search, either exhaustively (grid search) or by sampling search points at random (random search). As we show in Section 5, this simple approach can be very effective in practice. When some of the mutators are infinite, Sammo offers iterative search algorithms to optimize prompts. Algorithm 1 shows the skeleton of how Sammo implements iterative search. Starting with an initial baseline metaprompt ($\pi_{0}$) and small amount of training data as input, iteratively modifies a current set of candidates $\Pi_{\text{active}}$ through mutations to generate a new generation of candidates.
Specific choices for the functions InitializeCandidates, SampleCandidates, condition, and Prune in Algorithm 1 yield common search algorithms such as beam search, regularized evolutionary search Real et al. (2019) or breadth-first search. For example, beam search typically starts with one candidate and then selects all parents at the current depth, keeping only the top $k$ performing children for the next round. The novelty of Sammo lies not in this outer search but in the fact that we represent the higher level structure of a metaprompt explicitly which in turn allows us to define a rich set of operators that can both transform the structure as well the content.

### v2

# Symbolic Prompt Program Search: A Structure-Aware Approach to Efficient Compile-Time Prompt Optimization

@@ Abstract
In many modern LLM applications, such as retrieval augmented generation, prompts have become programs themselves. In these settings, prompt programs are repeatedly called with different user queries or data instances. A big practical challenge is optimizing such prompt programs. Recent work has mostly focused on either simple prompt programs or assumed that the general structure of a prompt program is fixed.
We introduce SAMMO, a framework to perform symbolic prompt program search for compile-time optimizations of prompt programs. SAMMO represents prompt programs on a symbolic level which allows for a rich set of transformations that can be searched over during optimization. We show that SAMMO generalizes previous methods and improves the performance of complex prompts on (1) instruction tuning, (2) RAG pipeline tuning, and (3) prompt compression, across several different LLMs. We make all code available open-source at https://anonymous.4open.science/r/sammo-4003.
@@ 1 Introduction
Figure 1: Left: Symbolic prompt program (SPP) for a binary classification task, where each node is a function with attributes and dependencies (children). The example also shows how the SPP allows for structural changes (e.g., DeleteNode) and attribute-based changes (e.g., ChangeFormat) which, after applying, result in the mutated prompt (Right). These enable Sammo to explore a large set of possible prompt candidates automatically.
@@
With the recent development of large language models (LLMs) such as Mixtral 8x7B Jiang et al. (2024) or GPT-4, it is now possible to provide LLMs with longer inputs including richer context and more detailed instructions. As a result, the complexity of prompts has increased, including longer strings of instructions, demonstrations or examples, and specification of more structured output. These lengthy instructions are often reused as part of larger prompts, where static information (e.g., instructions) is combined with input-dependent information (e.g., user queries, retrieved documents in RAG systems) at runtime. This has led to prompts being regarded as programs themselves, where each part is the result of other subprograms Khattab et al. (2023).
As LLM performance depends on a variety of factors such as paraphrasing, formatting and ordering Sclar et al. (2023); Liu et al. (2024), a key challenge to solve in practice is automatically finding the best prompt program for a given task and model. In this paper, we focus on compile-time optimization of prompt programs under black-box LLM access. Compile-time optimization is carried out only once before deployment and has two large advantages: (i) we can amortize the optimization cost over multiple uses of the prompt program and (ii) keep the existing run-time architecture unchanged. This is different from run-time optimization approaches, e.g., token pruning Jiang et al. (2023) which are called every time before inference and thus have non-amortizable costs.
@@
A lot of work on prompt optimization has focused on simple prompt programs where the program consists of a single string (which is subsumed by our representation in Section 2), mostly using structure-unaware operations such as paraphrasing Chen et al. (2023); Pryzant et al. (2023); Zhou et al. (2023). Another line of work has focused on optimizing prompts with more complex programmatic structure, but still considers only static prompt structures, e.g., DSpy Khattab et al. (2023). Initial work in this direction focused on applying textual mutators to different prompt components Fernando et al. (2023); Ye et al. (2023); Khattab et al. (2023). Subsequent work considered both textual mutation and hyperparameter selection Sclar et al. (2023). However, with increasing complexity of prompt structure, many prompt optimization techniques are no longer applicable and a new approach is needed that is able to optimize complex prompt programs with their hyperparameters.
To address this, we introduce Sammo, a general purpose framework for compile-time optimizations of prompt programs. Different from existing approaches Sammo represents prompt programs as symbolic prompt programs (SPP) which are abstract program graphs that can be changed arbitrarily. This allows Sammo to efficiently search through the space of valid prompt programs by automatically generating new promising candidates through mutations. Moreover, Sammo allows to both represent the internal structure of prompts (e.g., sections) as well as program structure (e.g., calling a retriever) in a unified fashion. By writing custom mutation operators, practitioners can encode their domain knowledge and guide the search process. Sammo naturally extends previous prompt programming approaches such as DSpy (Khattab et al., 2023) and encompasses several specialized prompt tuning methods as special cases. Through SPPs, Sammo (i) allows dynamic changes to prompt programs, (ii) generalizes compile-time optimizations to all prompt components (e,g., textual content, hyperparameters).
@@
Sammo is a general framework for prompt optimization in a black-box setting. We demonstrate the utility of symbolic prompt program search (SPPS) in Sammo in three scenarios: (1) instruction tuning, (2) retrieval augmented generation (RAG) pipeline tuning, and (3) prompt compression for annotation tasks. Our experiments show that Sammo generalizes previous methods and provides gains of 10-100% in instruction tuning, 26-133% in RAG tuning, and over 40% in prompt compression in performance. Moreover, we show that complex prompts need to be optimized separately for each LLM and that gains are more pronounced for weaker LLMs. Sammo code is available at https://anonymous.4open.science/r/sammo-4003 under an MIT license.
2 Symbolic Prompt Programs
A prompt program $\pi$ is a function that takes input data $X$ and maps it to another string $\hat{Y}=\pi[X]$. Sammo’s biggest innovation is the use of symbolic prompt programs to represent valid prompt structures, enabling flexible transformations and efficient search. That is in contrast to frameworks such as DSPy Khattab et al. (2023), that use static prompt programs in their optimization process.
To illustrate the advantages of symbolic prompt programs, consider the example prompt program in Figure 1 for a binary classification task. Each node is a function with attributes whose dependencies are indicated through child nodes. The program combines a number of text spans with the input data, sends it to the LLM and parses the output response. Static programs (e.g., DSPy programs) have a fixed structure which limits the operations to mostly changes in a node’s attributes, e.g., by paraphrasing text. Symbolic prompt programs (SPPs) that Sammo uses allow us to make much larger changes, all while ensuring that the prompt program remains semantically valid. Two useful operators are highlighted in Figure 1, where we change the data format to xml and also remove the instructions prefix. These program modification would be challenging, if not infeasible, with static programs. Section 2 will introduce many more possible operators for SPPs.
2.1 Graph Representation of SPPs
We represent prompt programs as directed acyclic graphs where nodes are functions and edges indicate call dependencies. Each prompt program $\pi[\cdot]$ is a DAG $G_{\pi}=(V,E)$ where nodes $v=(f_{v},\theta_{v})$ are functions (or subprograms) $f_{v}$ with attributes $\theta_{v}$. For example, in Figure 1, the left-most node is a function that renders its single attribute “Instructions:”. The symbolic part of the SPP describes the fact that we represent the entire $G_{\pi}$ symbolically in our implementation through pyGlove Peng et al. (2020).
2.2 Efficient Execution of SPPs
Ultimately, we need to compute $\pi[X]$ for different $\pi$ and $X$ during execution. To compute $\pi[X]$ on $G_{\pi}$, information will be passed top-down and then aggregated bottom-up. We start by calling the root node $\pi[X]=f_{v_{\mathrm{r}}}(\mathrm{state}=\{X\})$. For convenience, we divide the execution of every function $f_{v}$ into two non-recursive processing steps, i.e., $f_{v}=(f_{v}^{\mathrm{top-down}},f_{v}^{\mathrm{bottom-up}})$. Then, we evaluate each node $f_{v}$ recursively,
$\mathrm{state}^{\prime}\leftarrow f_{v}^{\mathrm{top-down}}(\mathrm{state},\theta_{v})$
$Y\leftarrow\big{\{}f_{u}(\mathrm{state}^{\prime})\mid(v,u)\in E\big{\}}$
$\textbf{return}{}f_{v}^{\mathrm{bottom-up}}(\mathrm{state}^{\prime}\cup\{Y\},\theta_{v}).$
For example, $v=\;$RenderInputExamples from Figure 1 will only execute a $f_{v}^{\mathrm{top-down}}$ function which formats the raw data into strings (using the format attribute).
3 Compile-time Prompt Optimization
Our goal in this paper is the compile-time optimization of prompt programs. Compile-time indicates that the optimization is done only once before using the prompt program, typically many times over with different inputs. We assume a black-box model for LLM access where the only information returned is the response text and no probabilities.
Given an objective $S_{\phi}(Y,\hat{Y})\in\mathbb{R}$ and set of samples $D_{s}$ drawn i.i.d from a data distribution $P(X,Y)$, the compile-time optimizer’s goal is to return a more performant prompt program:
$\hat{\pi}=\rho_{\text{compile}}(D_{\text{s}},S_{\phi}).$ (1)
Note that $S_{\phi}$ can also specify trade-offs between potentially multiple individual quality objectives $S_{i}(Y,\hat{Y})$. Equation 1 is different but complementary to run-time optimization which is run every time when input $X$ arrives (i.e., $\tilde{X}=\rho_{\mathrm{run}}(\pi[X]))$ and can tailor its decisions to $X$.
We implement $\rho_{\text{compile}}$ as a search procedure over a search space $\Pi$ of (symbolic) prompt programs. Our goal is to find a prompt program $\pi$ that performs best across the entire data distribution:
@@ 3 Compile-time Prompt Optimization
Here, $P(X,Y)$ is the distribution that the labeled samples $D_{s}$ are drawn from. Since the data distribution $P(X,Y)$ is unknown and the evaluation cost of Equation 2 is prohibitive, we resort to using the following sampled score in the spirit of empirical risk minimization (ERM):
@@
For the experiments in this paper, we assume that dataset sizes are on the order of hundreds of examples. This represents a reasonable amount of data that can be hand-labeled; increasing the amount would limit the applicability in practice substantially.
@@ 4 Sammo: Structure-Aware Multi-objective Metaprompt Optimization
Figure 2: Sammo is a flexible framework for structured prompt optimization, and offers two classes of search algorithms depending on the set of mutators used.
@@
11:     $\Pi_{\mathcal{C}}\!\leftarrow\!\!\textsc{Prune}(\Pi_{\mathcal{C}}\cup\Pi_{\text{nextgen}},D_{\text{train}},S_{\phi})$
@@
We outline Sammo, which is a general framework for optimizing the performance of symbolic prompt programs. Specific details can be found in the open-source implementation of Sammo. Sammo uses two classes of search algorithms with a rich set of mutation operators to explore the space of prompts.
The main novelty of Sammo is that prompt programs are represented with all their structural dependencies on a symbolic level. Without it, one is bound to a rigid prompt structure and also limited to a much smaller set of possible mutation operators. Moreover, it is also challenging to describe the search space efficiently. Sammo supports both explicit and implicit descriptions of the search space through symbolic operators.
Figure 2 presents an overview of Sammo’s main components. Starting from the top, Sammo employs two types of search algorithms for different scenarios. Below that sits of layer of prompt mutation operators that define the search space and actions. At the bottom is a base layer of function operators from which prompt programs are being composed, corresponding to the nodes in Figure 1.
4.1 Search Algorithms
A big challenge in practice is defining the right search space $\Pi$ for prompt optimization. To help with this, Sammo supports two different ways of specifying the search space which then correspond to a class of search algorithms.
Enumerative Search. First, similar to hyper-parameter search, in some instances the search space can be defined explicitly as a set of choices that are known a-priori. This leads to search algorithms that can leverage the resulting grid, a class of search strategies we call enumerative algorithms. As we show in Section 5.2, this simple approach can be very effective in practice. In its current version, Sammo implements grid search and random search as enumerative search strategies. In summary, enumerative search is useful when choices are known, the search space is small and to get a quick picture of the variability of different choices.
Iterative Search. Second, sometimes the search space can also be described implicitly through a starting state as well as a set of mutation operators that can be applied. This is useful in cases where the set of valid choices is not known a-priori, e.g., all valid paraphrases of a sentence. Sammo supports this specification through iterative search strategies. Sammo implements a generic template as a basis for several well-known algorithms and derives more concrete search algorithms from it. Algorithm 1 shows the skeleton of how Sammo implements iterative search. Starting with an initial prompt program ($\pi_{0}$), it iteratively modifies a current set of candidates $\Pi_{\text{active}}$ through mutations to generate a new generation of candidates.
Specific choices for the functions InitializeCandidates, SampleCandidates, condition, and Prune in Algorithm 1 yield common search algorithms such as beam search, regularized evolutionary search Real et al. (2019) or breadth-first search. We use beam search for all our experiments in this paper but will explore more sophisticated search strategies in future work.
@@ 4.2 Prompt Mutation Operators
Table 1: Examples for mutation operators, grouped by what part of a SPP $\pi$ they affect. Sammo allows for a rich set of operations whereas traditional prompt optimization techniques only focused on operations that change the text.
@@
At the heart of Sammo optimization are mutation operators. Formally, a mutation operator is a probabilistic function $m:\Pi\times\Pi\longrightarrow[0,1]$ that specifies how to transition from a SPP $\pi$ to an edited version $\pi^{\prime}\in\Pi$. This structure-aware component of Sammo opens up a new class of operators, for example operators that only modify specific sections or paragraphs. These can range from trivial (e.g., rephrasing a sentence) to complex (e.g., inducing a new set of task instructions).
Table 1 shows a non-comprehensive set of mutation operators, grouped by what part of an SPP they change. Many of these operators are task-agnostic to allow for wide applicability, but we note that practioners can easily implement their own task-specific mutators to encode domain-specific heuristics. To the best of our knowledge, Sammo is the first optimization method that can also optimize for large structural changes and data formatting.